### v1

# A neural network for estimating compact binary coalescence parameters of gravitational-wave events in real time

Abstract
Low-latency pipelines analyzing gravitational waves from compact binary coalescence events rely on matched filter techniques. Limitations in template banks and waveform modeling, as well as non-stationary detector noise cause errors in signal parameter recovery, especially for events with high chirp masses. We present a quantile regression neural network model that provides dynamic bounds on key parameters such as chirp mass, mass ratio, and total mass. We test the model on various synthetic datasets and real events from LIGO-Virgo-KAGRA gravitational-wave transient catalogs. We find that the model accuracy is consistently over 90% across all the datasets. We explore the possibility of employing the neural network bounds as priors in online parameter estimation. We find that they reduce by 9% the number of likelihood evaluations. This approach may shorten parameter estimation run times without affecting sky localizations.
- TPR
- True Positive Rate
- FPR
- False Positive Rate
- BNS
- binary neutron star
- NSBH
- neutron star-black hole
- MDC
- Mock Data Challenge
- EoS
- equation of state
- NS
- neutron star
- GW
- gravitational-wave
- GWs
- gravitational waves
- LIGO
- Laser Interferometer GW Observatory
- aLIGO
- Advanced LIGO
- AdVirgo
- Advanced Virgo
- O4
- fourth observing run
- LVK
- LIGO, Virgo, and KAGRA
- LV
- LIGO-Virgo
- O1
- first observing run
- O2
- second observing run
- O3
- third observing run
- oLIB
- Omicron+LIB
- KAGRA
- KAmioka GRAvitational–wave observatory
- GraceDB
- GRAvitational-wave Candidate Event DataBase
- LLAI
- Low-Latency Alert Infrastructure
- CBC
- compact binary coalescence
- BBH
- binary black hole
- FAR
- false alarm rate
- GstLAL
- GStreamer LIGO Scientific Collaboration Algorithm Library
- SVD
- singular value decomposition
- SNR
- signal–to–noise ratio
- FGMC
- Farr, Gair, Mandel and Cutler formalism of Poisson Mixture Model
- MBTA
- Multi-Band Template Analysis
- SPIIR
- Summed Parallel Infinite Impulse Response
- GPU
- Graphical Processing Units
- cWB
- Coherent WaveBurst
- GCN
- General Coordinates Network
- SCiMMA
- Scalable CyberInfrastructure for Multi-Messenger Astrophysics
- BAYESTAR
- BAYESian TriAngulation and Rapid localization
- FITS
- Flexible Image Transport System
- HEALPix
- Hierarchical Equal Area isoLatitude Pixelization
- MOC
- Multi-Order Coverage
- EM
- electromagnetic
- NS
- neutron star
- BH
- black hole
- LLOID
- Low Latency Online Inspiral Detection
- IIR
- Infinite Impulse Response
- ToO
- Target of Opportunity
- PE
- Parameter Estimation
- Bilby
- Bayesian Inference Library for CBC GW signal
- RAVEN
- Rapid, on-source VOEvent Coincident Monitor
- MMA
- Multi-Messenger Astronomy
- GWTC
- Gravitational-Wave Transient Catalog
- ML
- Machine Learning
- NN
- Neural Network
- QR
- Quantile Regression
- ROQ
- Reduced Order Quadrature
I Introduction
Over the course of three observing runs, LIGO, Virgo, and KAGRA (LVK) have confirmed about compact binary coalescence (CBC) detections of binary neutron stars, binary black holes and neutron star-black holes Abbott et al. (2019, 2021a, 2021b). The fourth observing run (O4) 111https://observing.docs.ligo.org/plan has seen more than 200 public alerts of gravitational-wave (GW) triggers 222https://gracedb.ligo.org/superevents/public/O4/.
The majority of these alerts are generated from CBC data analysis pipelines (Ewing et al., 2023; Dal Canton et al., 2021; Aubin et al., 2021; Chu et al., 2021) which are capable of detecting and analyzing triggers with latency less than 20 seconds Sharma Chaudhary et al. (2024). The critical component of this process is matched filter (Allen et al., 2012; Cannon et al., 2012). First, the detector data are filtered with a prebuilt template bank of waveforms Sakon et al. (2022). Then the pipelines provide point estimates of the signal parameters by selecting the template with the highest signal–to–noise ratio (SNR) or another suitable detection statistic Messick et al. (2017); Tsukada et al. (2023); Dal Canton et al. (2021); Aubin et al. (2021); Chu et al. (2021).
Various injection studies show that systematic and statistical errors affect these point estimates Ewing et al. (2023); Chatterjee et al. (2020). These uncertainties are especially relevant for high-mass BBH systems because of the smaller number of templates in that region of the bank Baird et al. (2013). Additional uncertainties arise from the need to reduce the size of the template banks Capano et al. (2016). Therefore, the parameters of CBC signals are not known in low latency with good accuracy; a full-scale Parameter Estimation (PE) on time scales of hours to days is required to extract their best estimates.
This situation is far from optimal. The physical parameters of BNS and NSBH mergers, in particular the chirp mass and total mass of the system, determine the disk wind and dynamic ejecta Margalit and Metzger (2019), and ultimately the post-merger evolution of the remnant. Therefore, estimating accurate parameters in low latency is critical to produce light curves and event sky localizations for electromagnetic (EM) follow-up observations.
Machine Learning (ML) methods Chatterjee et al. (2020); Berbel et al. (2024); Sharma Chaudhary et al. (2024) have been successful in identifying source properties in real time, such as the presence of a neutron star (NS), the possibility of a post-merger remnant, the presence of mass-gap and sub-solar mass compact objects. However, obtaining accurate point estimates of CBC parameters may prove challenging because of pipeline bias. In this paper, we take a different approach and develop a ML-based method to provide real-time bounds on chirp mass, mass-ratio, and total mass parameters of a CBC signal.
We train a Neural Network (NN) Quantile Regression (QR) Bassett et al. (1978) algorithm on a dense dataset of simulated CBC events. We derive bounds on chirp mass, mass-ratio, and total mass in the detector frame corresponding to selected quantile values. We find that the algorithm’s accuracy is over 90%. One of the key techniques that have been adopted to reduce the run time of online PE is limiting the prior space of the sampling algorithm Morisaki et al. (2023a); Rose et al. (2022a); CaitlinRose (2024). Our method could provide a dynamic way for doing so.
The paper is organized as follows. In Section II we describe the NN architecture and the datasets for training and testing. In Section III – VI, we present the results of the algorithm on various synthetic datasets and real events from the Gravitational-Wave Transient Catalog (GWTC). In section VII we explore the possibility of using the NN output to define priors for PE in low latency. Conclusions are presented in Section VIII.
II Data and model architecture
We consider two different datasets in our analysis. The first dataset has been utilized in the literature to train ML-based algorithms for source property inference Chatterjee et al. (2020); Sharma Chaudhary et al. (2024); Berbel et al. (2024). It contains approximately simulated CBC signals injected in detector noise from the LIGO-Virgo (LV) second observing run (O2) and filtered by the GstLAL search pipeline Ewing et al. (2023) with a false alarm rate (FAR) smaller than 1/month. The properties of this dataset are given in Table 1 of Ref. Chatterjee et al. (2020). We use this dataset for training, validation and testing purpose.
The second dataset is taken from the LVK third observing run (O3) replay Mock Data Challenge (MDC) Sharma Chaudhary et al. (2024). It consists of simulated CBC signals that were analyzed by online LVK search pipelines before the O4 in order to evaluate their performance. We utilize this dataset solely for testing purposes due to the limited number of simulated signals that the pipelines successfully recover. We also utilize a subset of the O3 replay MDC data for our PE analysis in subsection VII.
Both sets contain the intrinsic parameters of the simulated signals (detector frame masses and spins) and the template parameters that are recovered by the search pipelines:
As the chirp mass of the signal increases, the error in its pipeline recovered value also increases Ewing et al. (2023). Therefore, following a binning approach adopted by the online PE algorithm bilby-pipe 1.3.0
, we split the data into four different bins according to the value of the chirp mass. Table 1 lists the lower and upper bounds of these bins as well as their sizes. In order to train and test our method, we further split the datasets into () train, validation, and test sets.
Lower Bound | Upper Bound | O2 dataset | O3 replay MDC | |
BIN A | 23187 | 1049 | ||
BIN B | 9374 | 1049 | ||
BIN C | 120937 | 1325 | ||
BIN D | 49909 | 2082 |
Figures 11, 12, and 13 show the performance of the GStLAL pipeline on the O2 dataset (left panels) and the performance of the O4 pipelines on the replay MDC dataset (right panels) for different bins and different parameters. The maximum injected value for and in the MDC dataset is higher than in the O2 dataset. The template bank for matched filtering in the MDC also covers a larger region of the parameter space compared to O2. As a consequence, the range of the recovered parameters is larger for the MDC dataset than the O2 dataset. Since we train solely on O2 data, we restrict the values of the MDC parameters within the O2 bounds. This reduces the number of injections for the MDC dataset listed in Table 1 to 594, 604, 668, and 262 injections for bins A, B, C, and D, respectively.
Our NN model architecture is based on QR. Regression is a supervised ML technique that derives a relationship between a response variable and its input vector . Conventional regression utilizes the least squares method to give a point estimate of the response variable as a conditional mean over . QR is a variant of regression that estimates the conditional quantiles of the response variable.
Typically, NN implementations of QR are more accurate than conventional nonlinear QR methods Xu et al. (2017). The training of a QR NN model requires minimizing a quantile loss function. Given a random sample of response variables , where , the loss function for a quantile is Cannon (2011)
where is the NN estimate of the quantile and
In the case of a single NN model predicting multiple quantiles , where , the loss function can be generalized to
Our NN QR configuration is illustrated in Fig. 1 and described below following the PyTorch Ansel et al. (2024) library nomenclature Foundation (2025). The architecture consists of a fully connected NN with two (24, 12) hidden layers interleaved with dropout and LeakyReLU
activations. The input is pre-processed by a LayerNorm
layer. The output of the hidden layers is sorted in ascending order by a SoftSort
. This configuration prevents possible inversions of quantile values in the output Blondel et al. (2020). The latter is a set of unconstrained real numbers that is conditioned according to the specific training parameter: chirp mass, , mass ratio, , and total component mass . For the mass ratio we apply a sigmoid function to constrain the output of the sorting layer to the physical range . In the case of chirp mass and total component mass, we take the exponential of the sorting layer output and multiply it by the corresponding recovered parameters. For example, in the case of the chirp mass we have , where the prime indicates the output of the sorted layer.
We utilize the AdamW
optimizer, the CosineAnnealingWarmRestarts learning rate scheduler, and initial learning rates equal to () for and (). We train with the stochastic gradient descent and loop through the dataset in 100 batches of samples (epochs). The model weights are selected across epochs to minimize the loss on the validation set. The optimal training hyperparameters are selected by testing multiple configurations of the optimizer, learning rate, and the number of nodes in the hidden layers. The final choice is based on the configuration that minimizes the validation loss.
Figure 2 displays the loss curves for the twelve independent models (four chirp mass bins three parameters). As the number of the training epoch increases, both training and validation losses decrease, becoming essentially constant after 50 epochs. This is due to the annealing of the learning rate.
To assess the performance of our models we define two metrics, accuracy and interval width. For each quantile with target accuracy we build the confidence interval , where is any of the training parameters. If the true (injected) value of the parameter lies within the interval, we call it accurate. The fraction of accurate intervals across the dataset defines the accuracy of the models. The interval width is defined as . An optimal model produces intervals with maximum accuracy and smallest width compatible with the uncertainties of the recovered parameter.
III O2 results
Figure 3 shows the performance of the models on the O2 testing dataset. The dashed diagonal lines indicate ideal performance. The model for the mass ratio slightly underperforms for the A and C bins. This could be an effect of the final sigmoid layer for the model. The sigmoid function only asymptotically reaches and . Therefore, the NN output may be less accurate for equal or extreme mass ratio events.
Bin | Parameter | Accuracy | Median width |
0.916 | 0.512 | ||
A | 0.968 | 0.002 | |
0.974 | 0.234 | ||
0.962 | 0.642 | ||
B | 0.961 | 0.007 | |
0.975 | 0.977 | ||
0.916 | 0.893 | ||
C | 0.976 | 0.159 | |
0.970 | 17.217 | ||
0.961 | 0.731 | ||
D | 0.963 | 17.232 | |
0.966 | 40.789 |
Table 2 lists the accuracy and interval median width for different NN models and bins. The accuracy is estimated on the testing sets after setting a 96% target accuracy on the training set. The accuracy is consistently above 90% for all models and bins. Figure 4 shows that the NN interval widths are typically larger than the pipeline errors with their largest values being comparable. Reducing the target accuracy leads to smaller width distributions.
IV O3 MDC results
Bin | Parameter | Accuracy | Median width |
0.983 | 0.522 | ||
A | 0.966 | 0.003 | |
0.981 | 0.284 | ||
0.891 | 0.643 | ||
B | 0.947 | 0.005 | |
0.944 | 0.832 | ||
0.978 | 0.918 | ||
C | 0.964 | 0.145 | |
0.984 | 14.657 | ||
0.897 | 0.825 | ||
D | 0.969 | 19.83 | |
0.954 | 70.291 |
Figure 5 and Table 3 show the results of the NN models trained on O2 data on the O3 MDC set. Since the parameter space of the O2 and the O3 MDC datasets are different, we evaluate the models on a reduced O3 dataset (see Sec. II.) The different trends of the parameters in the bins wrt O2 are likely due to the inclusion of multiple pipelines in the O3 dataset, while training on the output of a single pipeline (GstLAL) in the O2 data. Changes in pipelines and their template banks leading up to O4 also affect injection recovery across different datasets. In spite of this, Fig. 5 and table 3 confirm that the NN models perform similarly on the O3 set as they do on O2.
Figure 6 illustrates the output of the chirp mass model for 96% target accuracy across three random injected chirp mass ranges. The three panels show the pipeline recovered chirp mass and the NN predicted intervals vs. the injected chirp mass for a sample of three different ranges. The pipelines recover the chirp mass fairly accurately between [1.25, 1.28] (top panel). Therefore, the NN predicted intervals are of the order of . Even when the injected chirp mass is not accurately recovered by the pipelines, the NN intervals still encompass the true injected value. As the injected increases, the pipeline accuracy decreases (middle panel). The median interval width also increase, ensuring that the injected remains included within the NN prediction. In the [30, 40] range (bottom panel), the pipelines underestimate most of the injected chirp masses. The NN model compensates for this bias by dynamically extending the interval upper bound. Overall, the NN provides accurate intervals for the O3 Replay MDC even if it was trained on O2 data.
V Performance on a synthetic dataset
In order to assess the NN performance on the full parameter space, we create a dataset of synthetic injections. We choose a log-uniform distribution for between 0.5 and 55 and a uniform distribution for between 0.05 and 1. We set the spin components to zero and fix the SNR to 10, 15, 20, and 25. Figure 7 shows the NN interval width with 96% target accuracy. Overall, in each bin the interval width decreases as increases. The width also generally increases with at fixed . These trends can also be seen in Fig. 8 which shows the widths of the chirp mass intervals for different values of the SNR. The median widths decrease with increasing SNR and other input parameters held fixed. This is more evident for bins B, C, and D.
VI Performance on GWTC events
We further assess the performance of the NN models on the events from the GWTC Abbott et al. (2024, 2021b) that passed the low-latency alert threshold. We choose the input parameters for the NN from low-latency templates with highest SNR. The GWTC catalog provides median values and 90% symmetric credible intervals for the source parameters. The , , and columns in Table 4 list median detector frame values for these parameters obtained from the GWTC posterior distributions. We compare these values to the width intervals with 96% target accuracy.
The NN intervals of all events include their corresponding GWTC median values with the exception of GW190521_030229. This event is reported in GWTC with a chirp mass of , the largest in the catalog. However, its low-latency chirp mass estimate is . Moreover, the training dataset is restricted to events with chirp mass values below . Therefore, it is not surprising that the NN model underestimates the actual value of the event.
Event_id | GWTC | Interval | Interval | Interval | |||
GW190408_181802 | 2.1 | [20.7, 35.0] | [0.27, 0.98] | [49.7, 78.2] | |||
GW190412 | 2.1 | [14.1, 21.0] | [0.25, 0.98] | [32.6, 48.2] | |||
GW190421_213856 | 2.1 | [34.3, 66.5] | [0.27, 0.98] | [88.9, 152.3] | |||
GW190425 | 2.1 | [1.4847, 1.4893] | [0.34, 0.99] | [3.240, 3.944] | |||
GW190503_185404 | 2.1 | [15.9, 44.0] | [0.12, 0.98] | [46.3, 116.6] | |||
GW190512_180714 | 2.1 | [15.3, 27.1] | [0.25, 0.98] | [38.2, 62.3] | |||
GW190513_205428 | 2.1 | [24.0, 42.8] | [0.25, 0.98] | [59.5, 102.0] | |||
GW190517_055101 | 2.1 | [30.0, 46.6] | [0.35, 0.98] | [73.7, 114.0] | |||
GW190519_153544 | 2.1 | [45.2, 78.4] | [0.27, 0.97] | [113.2, 190.3] | |||
GW190521_030229 | 2.1 | [27.2, 88.8] | [0.09, 0.98] | [67.9, 255.1] | |||
GW190521_074359 | 2.1 | [34.3, 58.0] | [0.26, 0.98] | [80.4, 126.3] | |||
GW190602_175927 | 2.1 | [27.7, 82.0] | [0.18, 0.98] | [81.4, 211.3] | |||
GW190630_185205 | 2.1 | [23.8, 42.9] | [0.25, 0.98] | [60.0, 100.8] | |||
GW190701_203306 | 2.1 | [35.1, 72.9] | [0.28, 0.98] | [93.9, 173.8] | |||
GW190706_222641 | 2.1 | [60.4, 98.2] | [0.33, 0.99] | [142.6, 221.1] | |||
GW190707_093326 | 2.1 | [9.72, 10.66] | [0.01, 0.99] | [19.1, 62.4] | |||
GW190720_000836 | 2.1 | [10.16, 11.15] | [0.00, 0.99] | [18.7, 71.3] | |||
GW190727_060333 | 2.1 | [28.5, 52.4] | [0.31, 0.98] | [71.3, 122.8] | |||
GW190728_064510 | 2.1 | [9.79, 10.29] | [0.01, 0.98] | [20.2, 65.0] | |||
GW190814 | 2.1 | [6.29, 6.52] | [0.05, 0.99] | [15.2, 40.0] | |||
GW190828_063405 | 2.1 | [27.3, 49.4] | [0.26, 0.98] | [67.9, 114.2] | |||
GW190828_065509 | 2.1 | [14.6, 26.0] | [0.24, 0.98] | [36.1, 63.3] | |||
GW190915_235702 | 2.1 | [24.0, 36.9] | [0.31, 0.98] | [58.1, 93.6] | |||
GW190924_021846 | 2.1 | [6.30, 6.60] | [0.04, 0.96] | [12.4, 44.0] | |||
GW190930_133541 | 2.1 | [9.58, 10.11] | [0.02, 0.97] | [19.0, 77.7] | |||
GW191105_143521 | 3 | [9.37, 10.21] | [0.00, 0.99] | [18.6, 68.5] | |||
GW191109_010717 | 3 | [44.7, 80.0] | [0.33, 0.98] | [110.9, 189.6] | |||
GW191129_134029 | 3 | [8.33, 8.77] | [0.01, 0.98] | [15.0, 48.9] | |||
GW191204_171526 | 3 | [9.53, 10.01] | [0.02, 0.97] | [17.8, 55.5] | |||
GW191215_223052 | 3 | [21.3, 37.7] | [0.21, 0.98] | [53.1, 92.3] | |||
GW191216_213338 | 3 | [8.81, 9.30] | [0.04, 0.97] | [17.8, 63.2] | |||
GW191222_033537 | 3 | [23.3, 64.2] | [0.18, 0.98] | [66.7, 167.0] | |||
GW200105_162426 | 3 | [3.607, 3.761] | [0.01, 0.99] | [8.7, 27.8] | |||
GW200112_155838 | 3 | [30.4, 50.7] | [0.26, 0.98] | [73.5, 110.9] | |||
GW200115_042309 | 3 | [2.542, 2.646] | [0.11, 0.51] | [5.1, 10.5] | |||
GW200128_022011 | 3 | [38.3, 62.0] | [0.35, 0.98] | [91.5, 142.5] | |||
GW200129_065458 | 3 | [30.0, 45.2] | [0.26, 0.98] | [67.0, 98.0] | |||
GW200208_130117 | 3 | [32.7, 53.6] | [0.36, 0.98] | [78.4, 127.1] | |||
GW200219_094415 | 3 | [23.7, 51.8] | [0.23, 0.98] | [61.7, 131.3] | |||
GW200224_222234 | 3 | [34.6, 57.0] | [0.28, 0.98] | [82.8, 129.8] | |||
GW200225_060421 | 3 | [15.3, 24.8] | [0.27, 0.98] | [36.5, 57.9] | |||
GW200302_015811 | 3 | [25.9, 46.4] | [0.30, 0.98] | [63.1, 106.8] | |||
GW200311_115853 | 3 | [24.6, 38.1] | [0.27, 0.98] | [57.5, 93.7] | |||
GW200316_215756 | 3 | [10.3, 11.2] | [0.01, 0.98] | [21.7, 85.6] |
VII NN model outputs in low-latency PE
Bilby Ashton et al. (2019), is a Bayesian inference algorithm that employs the nested sampling technique with the Dynesty library implementation Speagle (2020) to explore the full parameter space of masses and spins. Bilby’s configuration in O4 takes into account uncertainties in detector calibration and marginalizes the posterior probability distribution over them. A Reduced Order Quadrature (ROQ) technique Canizares et al. (2015); Smith et al. (2016); Morisaki et al. (2023b) is implemented to reduce the computational cost of likelihood evaluations. Bilby determines the lower and upper bounds of chirp mass and mass ratio priors by estimating the pipeline uncertainty in recovering the template’s chirp mass. The width of this priors affects the convergence of the algorithm and the speed of the PE process. Since the NN can provide accurate narrower intervals compared to Bilby’s default choices, it could be a viable alternative for the first stage of the PE process.
In order to test this, we select 34 BBH, NSBH, and BNS events each with the highest SNR in two weeks of the O3 Replay MDC. We assume the boundary between NS and black hole (BH) to be . We run our models with a target accuracy of 98.8% and 96% for the mass ratio and the chirp mass, respectively. The higher target accuracy for is justified by the fact that in the low-mass region of the parameter space is well estimated by the pipeline. Therefore, a larger interval for is required to maximize the range of component masses searched by PE.
We produce full posteriors with bilby 2.2.1
and bilby-pipe 1.3.0
and uniform mass priors with intervals computed by the NN. We also perform PE on the same set with the default low-latency bilby listed in Table 5. We employ the relative binning likelihood Krishna et al. (2023) since the low-latency ROQ method is incompatible with the dynamical chirp mass intervals provided by the NN.
Low-latency Bilby runs with pre-defined chirp mass intervals set by the ROQ method. Because of this restriction, the PE process cannot initiate unless the NN estimates lie in these pre-defined intervals.
Conditions | Prior bounds on | |
Lower | Upper | |
2 | + 0.01 | |
4 | + 0.1 | |
8 | ||
9.57 | 5.21 | 10.99 |
12 | 8.71 | 20.99 |
16 | 10.03 | 19.04 |
25 | 12.8 | 31.8 |
45 | 18.8 | 62.8 |
29.9 | 199.9 |
In the nested sampling algorithm, each likelihood iteration replaces the lowest likelihood sample with a higher likelihood sample. This process efficiently converges on high-probability regions of parameter space. Narrow priors around the injected value typically lead to a faster convergence of the algorithm. Although this may produce posterior distributions with limited support, the method is valuable in low-latency applications where greater speed is preferable.
Figure 9 compares the number of likelihood evaluations for Bilby’s default configuration and with the NN priors. The NN priors show a % median reduction in the number of likelihood evaluations.
As a final check, we compare the event sky maps produced with NN and default priors. The searched area event distribution and the corresponding p-p plot are shown in Fig. 10. Both NN and default prior are consistent.
VIII Conclusions
Low-latency PE of GW CBC signals requires as an input accurate source parameter priors. We have presented a new NN method based on QR that yields confidence bounds for the chirp mass, mass ratio, and total mass source parameters of candidate events identified in online LVK searches. The convergence of the PE process relies on the width of the priors; the narrower the prior interval, the faster the PE calculation converges. The NN QR provides accurate estimates for chirp mass, mass ratio, and total mass intervals that are comparable to the errors from the search pipelines in recovering these parameters. Therefore, our method provides dynamic intervals narrower than those currently in use that still enable accurate low-latency PE while speeding up the process. It could be used to provide confidence bounds of event intrinsic parameters in real time, rather than point estimates that may suffer from search pipeline bias.
We train the NN on data from the second LV observing run and test it on O2 data, as well as data from the O3 MDC dataset and real events from GWTC. The method performs consistently across all datasets with an accuracy of over 90%.
The algorithm performance does not affect the quality of the event sky localizations; the searched area and searched probability derived from the NN priors are consistent with the current LVK implementation.
In our analysis we applied the NN method to the Bilby PE pipeline. The LVK is currently employing a second pipeline for low-latency PE. RAPIDPE-RIFT Pankow et al. (2015) is an algorithm that utilizes a non-Markovian sampler. The pipeline fixes the CBC intrinsic parameters on a grid and marginalizes over the remaining parameters to speed up the process. The initial grid region is determined from the template parameters. Therefore, the NN could be utilized to constrain the parameter ranges in the initial grid. Since the number of points in RAPIDPE-RIFT’s grid is fixed, a NN implementation would not decrease the latency of the pipeline. However, narrower intervals combined with RAPIDPE-RIFT’s adaptive mesh refinement Rose et al. (2022b) would increase the precision of the PE.
Our method could be further improved by training it on O4 datasets with better coverage of the parameter space. This approach would help the model increase its parameter space support and keep up to date with pipeline updates. Another refinement could be to extend the NN input to multiple independent pipelines and parameters not restricted to the highest SNR template.
IX Acknowledgments
This work is based upon work supported by the LIGO Laboratory which is a major facility fully funded by the U.s. National Science Foundation. We are grateful for computational resources provided by the LIGO Laboratory and supported by the U.S. National Science Foundation Awards PHY-0757058 and PHY-0823459. We express our gratitude to the many colleagues from the LVK collaborations, without whom this work would not have been possible. In particular, we thank Daniel Wysocki, Carl-Johan Haster, and Soichiro Morisaki for their help and productive comments. G.P. thanks Giovanni Andrea Prodi for useful discussions and the supervision of his master thesis.
S.S.C. is partially supported by the U.S. National Science Foundation under awards PHY-2011334 and PHY-2308693. This research was initiated during G.P.’s stay at the Institute of Multi-messenger Astrophysics and Cosmology, Missouri University of Science and Technology, as part of the 2023 INFN-NSF/LIGO Summer Exchange Program. G.P. acknowledges support from the Italian INFN (Istituto Nazionale di Fisica Nucleare) with scholarships n. 25384/2023 and 25912/2023. M.C. is partially supported by the U.S. National Science Foundation under awards PHY-2011334, PHY-2219212 and PHY-2308693.
This manuscript has been assigned LIGO Document Control Center number P2500183.